OpenStudy (anonymous):

looking at the top of a building. the angle of elevation to the top of the building is 38 degrees... the engineer know the building is 300 ft tall... what is the distance from the engineer to the base of the building to the nearest whole foot?

OpenStudy (anonymous):

tan 38 = 300/x (x= distance) so, x=300/tan38 use calculator & find the value... :)

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

welcome :)